OpenStudy (anonymous):

A high speed bullet train traveled 900 feet in 5 seconds. At this rate, how far would it travel in one minute?

OpenStudy (acxbox22):

\[\frac{ 900 }{ 5 }=\frac{ x }{ 60 }\] cross multiply \[5x=900(60)\] \[5x=54000\] x=10800 so the train will travel 10800 ft in 1 minute

OpenStudy (anonymous):

Hmm ur smart and ty!

OpenStudy (anonymous):

Malorie can work 10 math problems in 3 minutes. At this rate, how many problems can she work in one hour?

OpenStudy (anonymous):

Omg..